Find k so that the line kx-3y=4k is parallel to the line 2x-4y=1
Answer by dan_math(11)

If 2 lines are parallel then they have equal slope.
Rewrite your equations in a standard form: y=ax+b
kx-3y=4k;
-3y=-kx+4k;
y=(k/3)x+4k/3;
the slope of this equation is m1=k/3
the other equation:
2x-4y=1
-4y=-2x+1
y=(1/2)x-1/4
the slope of this equation is m2=1/2
since the lines are parallel, m1=m2 so k/3=1/2, k=3/2